Circular No. 6396

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

MARKARIAN 421 T. Weekes, Harvard-Smithsonian Center for Astrophysics, reports for

the Whipple Observatory Gamma Ray Collaboration on an extraordinary outburst of γ-ray activity from this previously detected BL-Lac object: “The source was observed on May 7 with the 10-m optical reflector under perfect observing conditions. Using the standard Whipple observing and analysis technique, we detected a signal at the 56σ level in 2 hr of observation with a doubling time of 1 hr. The energy threshold for the analysis was ∼ 350 GeV. Preliminary analysis gives the following counts (γ-rays/min): May 7.150–7.170 UT, 5.54 ± 0.27 (20.8σ); 7.180–7.199, 8.29 ± 0.33 (25.1σ); 7.200–7.219, 10.78 ± 0.36 (30.3σ); 7.220–7.240, 12.68 ± 0.40 (37.7σ); 7.240–7.244, 12.11 ± 0.91 (13.4σ). The average rate of 9.5 γ-rays/min is five times larger than for any previous observation (six times the level of the Crab Nebula) and is the brightest TeV signal recorded by us from any source to date. Subsequent observations with the Whipple telescope show that the source has declined to more usual levels (May 8, 0.62 ± 0.07 γ-rays/minute; May 9, 0.42 ± 0.09; May 10, 0.79 ± 0.10). Observations by P. Garnavich and N. Caldwell with the Whipple 1.2-m optical telescope indicate a brightening of 0.25 mag in the R band on May 7–10 compared with levels recorded the previous month. Markarian 421 was monitored in x-rays by XTE during May 3-9.”

COMET C/1996 B2 (HYAKUTAKE) C. G. Mason, R. D. Gehrz, T. J. Jones, and D. M. Williams, Uni-

versity of Minnesota (UM), report infrared photometry of comet C/1996 B2 using the UM/University of California at San Diego 1.52-m telescope (+ UM bolometer) at the Mt. Lemmon Observing Facility. Photometric magnitudes are averaged between Apr. 22.1 and 23.1 UT using a 13′′.5 circular aperture with a north-south throw of 129′′ between the source and reference beams: J = 4.70 ± 0.08, H = 4.26 ± 0.08, K = 3.17 ± 0.10, L = 0.43 ± 0.06, M = −0.96 ± 0.04, [7.8 μm] = −2.63 ± 0.03, [8.7 μm] = −3.29 ± 0.03, [9.8 μm] = −3.85 ± 0.03, [10.3 μm] = −4.11 ± 0.04, [11.6 μm] = −4.14 ± 0.05, [12.5 μm] = −3.69 ± 0.05, and [18.0 μm] = −4.29 ± 0.04. The infrared spectral energy distribution shows a dust temperature ∼ 40 percent higher than would be expected for small black spheres at the same heliocentric distance. The 10- and 20-μm silicate emission features are quite strong.

1996 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green